Issued $1,190,000 of five-year, 10% callable bonds dated July 1, Year 1, at a market (effective) rate of 11%, receiving cash of $1,145,149. Interest is payable semiannually on December 31 and June 30. Oct. 1. Borrowed $310,000 by issuing a 10-year, 7% installment note to Intexicon Bank. The note requires annual payments of $44,137, with the first payment occurring on September 30, Year 2. Dec. 31. Accrued $5,425 of interest on the installment note. The interest is payable on the date of the next installment note payment. Dec. 31. Paid the semiannual interest on the bonds. The bond discount amortization of $4,485 is combined with the semiannual interest payment. Year 2 June 30. Paid the semiannual interest on the bonds. The bond discount amortization of $4,485 is combined with the semiannual interest payment. Sept. 30. Paid the annual payment on the note, which consisted of interest of $21,700 and principal of $22,437. Dec. 31. Accrued $5,032 of interest on the installment note. The interest is payable on the date of the next installment note payment. Dec. 31. Paid the semiannual interest on the bonds. The bond discount amortization of $4,485 is combined with the semiannual interest payment. Year 3 June 30. Recorded the redemption of the bonds, which were called at 98. The balance in the bond discount account is $26,911 after payment of interest and amortization of discount have been recorded. (Record the redemption only.) Sept. 30. Paid the second annual payment on the note, which consisted of interest of $20,129 and principal of $24,008. Required: 1. Journalize the entries to record the foregoing transactions. For compound transactions, if an amount box does not require an entry,
Answer #1

Journal entries for the given transaction

date

entry

amount Dr ($)

Amount ($) CR

year1

(july 1)

cash A/c   

Discount on Bonds Payable A/c

To Bonds Payable

1145149

44851

1190000

( Being 10% Bonds issued on discount )

oct 1

Cash A/c

To Notes Payable A/c

310000

310000

( Being amount borrowed from bank for 10 years)

31 dec

Interest Expense A/c

To Interest Payable A/c

5425

  

5425

( Being Interest Accured on installment note )

31 dec

Interest Expense A/c

To discount on bonds Payable A/c

To cash A/c

63985

4485

59500

( being interest paid on bonds )

year 2

june 30

Interest Expense A/c

To discount on bonds Payable A/c

To cash A/c

63985

4485

59500

( being interest paid on bonds )

sep 30

interest expense A/c

interest payable A/c

notes payable A/c

To cash A/c

16275

5425

22437

44137

( Being Installment Paid on borrowed Money)

31dec

interest Expense A/c

To Interest Payable A/c

5032

5032

( Being Interest Accured on installment note )

31 dec

Interest Expense A/c

To discount on bonds Payable A/c

To cash A/c

63985

4485

59500

( being interest paid on Bonds)

Year 3

june 30

Bonds Payable A/c

Loss on Redemption of Bonds A/c

To Discounts on Bonds Payable A/c

To Cash A/c

1190000

7596

31396

1166200

( Being bonds are redeemed @98)

sep 30

Interest Expense A/c

Interest Payable A/c

Notes Payable A/c

To Cash

15097*

5032

24008

44137

( Being second installment of bank paid

* Total interest paid =20129

Interest payable which was accured = 5032

so interest expense = 20129-5032

=15097

Calculation of bonds Redemption

as discount on bonds =44851

so semi annual discount amortised =4485

total discount amortized till last day of redemption= 4485*3= 13455

amount left to amortised = 44851-13455

=31396

Bonds redmeption value = 1190000*98%

=1166200

discount on bonds left to be amortized = 31396

bonds payable = 1190000

so loss on redemption = 1190000-31396-1166200

= 7596